# Images Speak Louder than Words: Understanding and Mitigating Bias in Vision-Language Model from a Causal Mediation Perspective

### v1

@@ Abstract
Vision-language models (VLMs) pre-trained on extensive datasets can inadvertently learn biases by correlating gender information with specific objects or scenarios. Current methods, which focus on modifying inputs and monitoring changes in the model’s output probability scores, often struggle to comprehensively understand bias from the perspective of model components. We propose a framework that incorporates causal mediation analysis to measure and map the pathways of bias generation and propagation within VLMs. This approach allows us to identify the direct effects of interventions on model bias and the indirect effects of interventions on bias mediated through different model components. Our results show that image features are the primary contributors to bias, with significantly higher impacts than text features, specifically accounting for 32.57% and 12.63% of the bias in the MSCOCO and PASCAL-SENTENCE datasets, respectively. Notably, the image encoder’s contribution surpasses that of the text encoder and the deep fusion encoder. Further experimentation confirms that contributions from both language and vision modalities are aligned and non-conflicting. Consequently, focusing on blurring gender representations within the image encoder which contributes most to the model bias, reduces bias efficiently by 22.03% and 9.04% in the MSCOCO and PASCAL-SENTENCE datasets, respectively, with minimal performance loss or increased computational demands.
@@
Zhaotian Weng1, Zijun Gao1, Jerone Andrews2, Jieyu Zhao1 1University of Southern California, 2Sony AI wengzhao@usc.edu, zijungao@marshall.usc.edu, Jerone.Andrews@sony.com, jieyuz@usc.edu
@@ 1 Introduction
In this work, we propose a standardized framework to measure bias in VLMs, providing a comprehensive understanding of how bias flows within the entire model structure. Specifically, we use the GLIP model Li et al. (2022) as a case study, focusing on gender bias in the task of object detection, which is a predominant and challenging problem in computer vision. We conduct the analysis on both the MS-COCO Lin et al. (2014) and PASCAL-SENTENCE Rashtchian et al. (2010) datasets. We observe that GLIP model exhibits unbalanced inference capabilities on different genders, with certain indoor object categories like pets more likely to be associated with females and outdoor objects like vehicles with males. To holistically understand how the bias flows in the model, we adapt causal mediation analysis  Vig et al. (2020) to VLMs, providing a finer-grained study of contributions from different model components. We find that, among the different model components (text module, image module, and fusion module that combines them), the image module contributes the most to the model’s bias – over twice as much as the text module. In the MSCOCO and PASCAL-SENTENCE datasets, image features accounted for 32.57% and 12.63% of the bias generated, respectively, compared to approximately 15.48% and 5.64% by text features. Also, the interaction and updating process between image and text features during the deep fusion process significantly impacts bias production, accounting for about 57% of the contributions in the image and text encoders. Furthermore, by integrating interventions across different modules, we discovered that their contributions to bias are aligned and do not conflict, allowing us to prioritize bias mitigation efforts within the image encoder, which is the most substantial contributor to bias. Based on the results, we propose to effectively mitigate the bias in VLMs: reducing biases from the image module can successfully reduce bias by 22.03% on the MSCOCO dataset and 9.04% on the PASCAL-SENTENCE dataset, compared to a reduction of 7.08% and 1.18% in the text module. In summary, the contributions of our work are:
@@ 3.2 Causal Mediation Analysis Method
Figure 1: Causal Mediation Analysis example
@@ 4 Experimental Setup of Bias Measurement and Understanding
For the object detection task, we employed the GLIP model pre-trained on the O365, GoldG, CC3M, and SBU datasets Li et al. (2022). The model consists of an image module, a text module, and a deep-fusion module that updates and aligns image features and text features. For object detection, the GLIP model makes predictions based on the given image and a text input, which is a list of possible categories separated by commas.
@@
Initially, we implement replace-gender and mask-gender interventions on the inputs respectively without any alterations to the model components. By monitoring the changes in the values of BiasVL, the individual impacts of image and text inputs on gender bias within the input module were assessed. Subsequently, we conducted a detailed causal mediation analysis on the text encoder and image encoder, respectively, by choosing the attention head within a specific layer and those in all preceding layers as mediators, conducting experiments from shallow to deep layers. This analysis aimed to identify whether the text encoder or image encoder contributes more significantly to gender bias and to determine which layers in the model are principally responsible for bias generation. It also sought to understand how bias flows and accumulates across different layers within the encoders. Then, we selected a combination of attention layers from both the image encoder and text encoder as mediators to observe changes in bias and compare these results with previous findings, exploring whether different modalities reinforce bias or conflict in the direction of bias.
@@ 5.2 Bias Understanding with Causal Mediation Analysis
We conduct the causal mediation analysis on different modules to study their effect on the model bias. We find that the image module influences the model bias more than the text module and the fusion module. In addition, we show that the bias in the image and text modules are aligned – they are showing similar gender bias tendencies rather than conflicting ones.
@@
Applying the mask-gender intervention to the input image module reduced the BiasVL to 0.967 for the MSCOCO dataset and to 0.664 for the PASCAL-SENTENCE dataset, representing reductions of approximately 32.57% and 12.63%, respectively. We employed the attention heads in the image encoder as the mediator to examine both the indirect effects of this model component and the direct effects of the mask-gender on predictions. Figure 44(a) and Figure 44(e) illustrate that employing more attention heads as mediators leads to greater reductions in indirect effect, with diminishing reductions in direct effect. This supports an intuition that removing gender information from more layers in the image encoder weakens the model’s dependency on latent correlations between gender in images and specific objects, thus mitigating gender bias in predictions. Furthermore, while interventions at the input level significantly impact final predictions, targeting the image encoder alone achieves about 53% of the mask-gender effect.
@@
To further validate whether image features contribute more to bias creation than text features, we utilized the attention heads in the deep fusion encoder as the mediator, adjusting the attention heads’ parameters in the states of either mask-gender intervention or replace-gender intervention. The results displayed in Figure 44(d) and Figure 44(c) show that for the MSCOCO dataset, the indirect effects from mask-gender and replace-gender through the deep fusion encoder are up to 0.260 and 0.189, respectively, reducing the BiasVL by approximately 18.13% and 13.18%. For the PASCAL-Sentence dataset, the reductions are 10.80% and 0.53%, respectively (Figure 44(h) and Figure 44(g)). These findings reaffirm our conclusion that image features play a more substantial role in bias generation than text features. They also suggest that even though the deep fusion module does not extract features directly from images and text, the interactive updating process between text and image features significantly influences bias generation, accounting for approximately 55.70% of the effect observed with the encoder alone.
@@
MSCOCO PASCAL-S MSCOCO PASCAL-S MSCOCO PASCAL-S
GLIP-T 46.6 68.4 1.434 0.763 0 0
GLIP_ImageFair 46.2 68.3 1.118 0.694 22.03% 9.04%
GLIP_TextFair 46.6 68.4 1.322 0.754 7.8% 1.18%
Table 2: Performance comparison of different methods on MSCOCO and PASCAL-S (PASCAL-SENTENCE) datasets. AP (Average Precision) is the metric used for zero-shot object detection. “GLIP” represents the original GLIP model, “GLIP_ImageFair” denotes the model with bias mitigation implemented in the image encoder, and “GLIP_TextFair” refers to the model with bias mitigation applied in the text encoder. Intervention in the image encoder is more effective than the text encoder in reducing the bias score without significant performance loss.
@@ 7 Experimental Setup of Bias Mitigation
We evaluated the effectiveness of bias mitigation on the MSCOCO and PASCAL-SENTENCE datasets. To assess the model’s object detection performance, we compared it with the original GLIP  Li et al. (2022) on the MSCOCO and PASCAL-SENTENCE datasets using the AP (Average Precision) metric for zero-shot object detection.
@@ 9 Conclusion
Vision-language models (VLMs) trained on large-scale image-text pair corpora are at risk of learning social biases from their training data. In this paper, we introduced a standardized framework incorporating causal mediation analysis to measure and understand the pathways through which model bias is generated and propagated within VLMs. We discovered that image features contribute significantly more to model bias than text features, and the contributions from the image encoder substantially exceed those from the text encoder and deep fusion encoder. Furthermore, the contributions to bias from different language modalities reinforce each other. Subsequently, by focusing on the components that contribute most to bias, we efficiently reduced model bias.
Our work provides a framework for measuring, understanding, and mitigating model bias, which, although utilized here within the realm of object detection, can be extended to a wide range of VLM tasks. However, our framework is primarily applicable to white-box models, as it requires interventions at the internal components of the model. A promising direction for future work would involve expanding our framework to encompass additional modalities such as audio or video. This expansion could further enhance our understanding of multimodal interactions and their impact on bias, as well as deepen insights into how different sensory inputs contribute to, or mitigate, biases in AI systems.

### v2

@@ Abstract
Vision-language models (VLMs) pre-trained on extensive datasets can inadvertently learn biases by correlating gender information with specific objects or scenarios. Current methods, which focus on modifying inputs and monitoring changes in the model’s output probability scores, often struggle to comprehensively understand bias from the perspective of model components. We propose a framework that incorporates causal mediation analysis to measure and map the pathways of bias generation and propagation within VLMs. Our framework is applicable to a wide range of vision-language and multimodal tasks. In this work, we apply it to the object detection task and implement it on the GLIP model. This approach allows us to identify the direct effects of interventions on model bias and the indirect effects of interventions on bias mediated through different model components. Our results show that image features are the primary contributors to bias, with significantly higher impacts than text features, specifically accounting for 32.57% and 12.63% of the bias in the MSCOCO and PASCAL-SENTENCE datasets, respectively. Notably, the image encoder’s contribution surpasses that of the text encoder and the deep fusion encoder. Further experimentation confirms that contributions from both language and vision modalities are aligned and non-conflicting. Consequently, focusing on blurring gender representations within the image encoder which contributes most to the model bias, reduces bias efficiently by 22.03% and 9.04% in the MSCOCO and PASCAL-SENTENCE datasets, respectively, with minimal performance loss or increased computational demands.
@@ 1 Introduction
In this work, we propose a standardized framework to measure bias in VLMs, providing a comprehensive understanding of how bias flows within the entire model structure. Specifically, we use the GLIP model Li et al. (2022) as a case study, focusing on gender bias in the task of object detection, which is a predominant and challenging problem in computer vision. We conduct the analysis on both the MS-COCO Lin et al. (2014) and PASCAL-SENTENCE Rashtchian et al. (2010) datasets. We observe that GLIP model exhibits unbalanced inference capabilities on different genders, with certain indoor object categories like pets more likely to be associated with females and outdoor objects like vehicles with males. To holistically understand how the bias flows in the model, we adapt causal mediation analysis  Vig et al. (2020) to VLMs, providing a finer-grained study of contributions from different model components. We find that, among the different model components (text module, image module, and fusion module that combines them), the image module contributes the most to the model’s bias – over twice as much as the text module. In the MSCOCO and PASCAL-SENTENCE datasets, image features accounted for 32.57% and 12.63% of the bias generated, respectively, compared to approximately 15.48% and 5.64% by text features. Also, the interaction and updating process between image and text features during the deep fusion process significantly impacts bias production, accounting for about 57% of the contributions in the image and text encoders.
Furthermore, by integrating interventions across different modules, we discovered that their contributions to bias are aligned and do not conflict, allowing us to prioritize bias mitigation efforts within the image encoder, which is the most substantial contributor to bias. Based on the results, we propose to effectively mitigate the bias in VLMs: reducing biases from the image module can successfully reduce bias by 22.03% on the MSCOCO dataset and 9.04% on the PASCAL-SENTENCE dataset, compared to a reduction of 7.08% and 1.18% in the text module. In summary, the contributions of our work are:
@@ 3.2 Causal Mediation Analysis Method
Figure 1: Causal Mediation Analysis example. In this example, strength training serves as the treatment, aiming to improve athletic performance, while muscle relaxation acts as the mediator that indirectly affects the athletic performance. In our study, interventions in the input module directly affect gender bias in model predictions, while model components such as specific layers or neurons can mediate this effect.
@@ 4 Experimental Setup of Bias Measurement and Understanding
Figure 3: GLIP object detection pipeline. The input to the GLIP model consists of an image and a caption, which includes a list of possible categories separated by commas.
@@
For the object detection task, we employed the GLIP model pre-trained on the O365, GoldG, CC3M, and SBU datasets Li et al. (2022). The model consists of an image module, a text module, and a deep-fusion module that updates and aligns image features and text features. For object detection, the GLIP model makes predictions based on the given image and a text input, which is a list of possible categories separated by commas. This input format, as highlighted by  Li et al. (2022), has demonstrated good performance. For example, in the case of the COCO dataset, the text input along with an image will be a text prompt like "teddy_bear . handbag . fork …… . baseball_glove . skateboard .", as illustrated in Figure 3.
@@
Initially, we implement replace-gender and mask-gender interventions on the inputs respectively without any alterations to the model components. By monitoring the changes in the values of BiasVL, the individual impacts of image and text inputs on gender bias within the input module were assessed. Subsequently, we conducted a detailed causal mediation analysis on the text encoder and image encoder. Previous studies  Vaswani (2017); Vig and Belinkov (2019) have emphasized the critical roles that attention heads and layers play in deep learning models. Therefore, we selected attention heads from these two modules as mediators. Specifically, we chose the attention heads within a specific layer, along with those in all preceding layers as mediators, and conducted experiments from shallow to deep layers. This analysis aimed to identify whether the text encoder or image encoder contributes more significantly to gender bias and to determine which layers in the model are principally responsible for bias generation. It also sought to understand how bias flows and accumulates across different layers within the encoders. Then, we selected a combination of attention layers from both the image encoder and text encoder as mediators to observe changes in bias and compare these results with previous findings, exploring whether different modalities reinforce bias or conflict in the direction of bias.
@@ 5.2 Bias Understanding with Causal Mediation Analysis
We conduct the causal mediation analysis on different modules to study their effect on the model bias. We find that the image module influences the model bias more than the text module and the fusion module. In addition, we show that the bias in the image and text modules are aligned – they are showing similar gender bias tendencies rather than conflicting ones. At this stage, we used four Quadro RTX 6000 GPUs. Since we primarily focus on inference tasks rather than training or fine-tuning, the computational resource requirements are not intensive.
@@
Applying the mask-gender intervention to the input image module reduced the BiasVL to 0.967 for the MSCOCO dataset and to 0.664 for the PASCAL-SENTENCE dataset, representing reductions of approximately 32.57% and 12.63%, respectively. We employed the attention heads in the image encoder as the mediator
to examine both the indirect effects of this model component and the direct effects of the mask-gender on predictions. Figure 55(a) and Figure 55(e) illustrate that employing more attention heads as mediators leads to an increase in the value of the indirect effect, while the direct effect diminishes. This supports an intuition that removing gender information from more layers in the image encoder weakens the model’s dependency on latent correlations between gender in images and specific objects, thus mitigating gender bias in predictions. Furthermore, while interventions at the input level significantly impact final predictions, targeting the image encoder alone achieves about 53% of the mask-gender effect.
@@
To further validate whether image features contribute more to bias generation than text features, we utilized the attention heads in the deep fusion encoder as the mediator, adjusting the attention heads’ parameters in the states of either mask-gender intervention or replace-gender intervention. The results displayed in Figure 55(d) and Figure 55(c) show that for the MSCOCO dataset, the indirect effects from mask-gender and replace-gender through the deep fusion encoder are up to 0.260 and 0.189, respectively, reducing the BiasVL by approximately 18.13% and 13.18%. For the PASCAL-Sentence dataset, the reductions are 10.80% and 0.53%, respectively (Figure 55(h) and Figure 55(g)). These findings reaffirm our conclusion that image features play a more substantial role in bias generation than text features. They also suggest that even though the deep fusion module does not extract features directly from images and text, the interactive updating process between text and image features significantly influences bias generation, accounting for approximately 55.70% of the effect observed with the encoder alone.
@@
MSCOCO LVIS MSCOCO PASCAL-S MSCOCO PASCAL-S
GLIP 46.6 17.6 1.434 0.763 0 0
GLIP_ImageFair 46.2 17.3 1.118 0.694 22.03% 9.04%
GLIP_TextFair 46.6 17.6 1.322 0.754 7.8% 1.18%
Table 2: Performance comparison of different methods. AP (Average Precision) is the metric used for zero-shot object detection. “GLIP” represents the original GLIP model, “GLIP_ImageFair” denotes the model with bias mitigation implemented in the image encoder, and “GLIP_TextFair” refers to the model with bias mitigation applied in the text encoder. Intervention in the image encoder is more effective than the text encoder in reducing the bias score without significant performance loss.
@@ 7 Experimental Setup of Bias Mitigation
We evaluated the effectiveness of bias mitigation on the MSCOCO and PASCAL-SENTENCE datasets. To assess the model’s object detection performance, we compared it with the original GLIP  Li et al. (2022) on the MSCOCO and LVIS datasets using the AP (Average Precision) metric for zero-shot object detection.
@@ 9 Conclusion
Vision-language models (VLMs) trained on large-scale image-text pair corpora are at risk of learning social biases from their training data. In this paper, we introduced a standardized framework incorporating causal mediation analysis to measure and understand the pathways through which model bias is generated and propagated within VLMs. We discovered that image features contribute significantly more to model bias than text features, and the contributions from the image encoder substantially exceed those from the text encoder and deep fusion encoder. Furthermore, the contributions to bias from different modalities reinforce each other. Subsequently, by focusing on the components that contribute most to bias, we efficiently reduced model bias.
Our work provides a framework for measuring, understanding, and mitigating model bias, which, although utilized here within the realm of object detection, can be extended to a wide range of VLM tasks. We used gender bias as a case study to showcase our methodology, as many influential studies in the natural language processing and fairness fields have begun with the study of gender bias   Bolukbasi et al. (2016); Zhao et al. (2019, 2017); Zhang et al. (2022b). Gender bias is one of the most representative types of bias, and understanding it can provide valuable insights for studying other forms of bias. The frameworks and evaluation metrics we developed for gender bias are generalizable and can be extended to other biases. For instance, if ground truth labels about age are available in the image data, researchers can modify the text descriptions to include age-related terms and apply the same methodology to study age-related bias. However, our framework is primarily applicable to white-box models, as it requires interventions at the internal components of the model. A promising direction for future work would involve expanding our framework to encompass additional modalities such as audio or video. This expansion could further enhance our understanding of multimodal interactions and their impact on bias, as well as deepen insights into how different sensory inputs contribute to, or mitigate, biases in AI systems.